OpenStudy (anonymous):

2x+3/5 = x-2/6 solve for x .

OpenStudy (anonymous):

cross multiply : 12x+18=5x-10 7x=28 x=4

OpenStudy (anonymous):

12x + 18 = 5x -10 7x = -28 x = -4

OpenStudy (anonymous):

it is x=-4 ishaan ur right